OpenStudy (anonymous):

Find the derivative of the function. dy/dx= y=(3x^2-6x+1)^3

OpenStudy (ranga):

Use the power rule first and then the chain rule or Let z = 3x^2-6x+1 Then y = z^3 dy/dx = (dy/dz)(dz/dx)

jigglypuff314 (jigglypuff314):

use the power rule which is: y = x^n then y' = n x^(n-1) and chain rule... y = (stuff)^3 then y' = 3(stuff)^(3-1) * deriv (stuff)

jigglypuff314 (jigglypuff314):

so y=(3x^2-6x+1)^3 y' = 3(3x^2-6x+1)^2 * deriv (3x^2-6x+1) = 3(3x^2-6x+1)^2 * (6x - 6) = 18(x - 1)(3x^2-6x+1)^2